1 The first section

We’ll use the following in Theorem 1.

a = b  (1)

Theorem 1. Equation (1) has nothing to do with (2) or (3). See also Lemma 2.

a^{2} = b^{2} + c^{2}  (2)
a = \sqrt{b^{2} + c^{2} − 2bc cos A}  (3)

Having had this pointless theorem, why not try for a lemma of similar class.

2 Another section

Lemma 2. Theorem 1 (in Section 1) contains two equations. While both are simple, (3) is more complex than (2).

3 The final section

This document includes Sections 1, 2 and 3. Lemma 2 is on page 1.